OpenStudy (anonymous):

(2,2)(-2,-1) is the slope 3/4 ?

OpenStudy (anonymous):

(x1,y2)=(2,2) (x2,y2)=(-2,-1) so y2=-1, y1=2 x2=-2, x1=2 slope= y2-y1/x2-x1 slope=(-1)-(2)/(-2)-(2) slope= -1-2/-2-2 slope=-3/-4 simplify slope=+3/4 answer? Yes!! more slope problems worked out step by step in math911 www.math911.com